two cyclists start from the same point and ride in opposite directions. One cyclist rides twice as fast as the other. In 3 h, they are 81 mi apart. Find the rate of each cyclist.
Answer by nerdybill(7384)

two cyclists start from the same point and ride in opposite directions. One cyclist rides twice as fast as the other. In 3 h, they are 81 mi apart. Find the rate of each cyclist.
.
You will be applying the distance formula:
d = rt
where
d is distance
r is rate or speed
t is time
.
Let x = speed of slower cyclist
then
2x = speed of faster cyclist
.
3(x + 2x) = 81
(x + 2x) = 27
3X = 27
X = 9 mph (slower cyclist)
.
faster cyclist:
2x = 2(9) = 18 mph (faster cyclist)